OpenStudy (anonymous):

HELP ME WILL FAN AND MEDAL!!! Julia is allowed to watch no more than 5 hours a week.(man that must suck) So far she has watched 1.5 hours. Write and solve and equality to show how many hours Julia can still watch this week?

OpenStudy (anonymous):

\[5-1.5\] should do it`

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

That's wrong... Lol...

OpenStudy (anonymous):

It's not an inequality.

OpenStudy (anonymous):

@Princess_Maria Ok, so she can't watch more than 5 hours a week, so which inequality symbol would it be?

OpenStudy (anonymous):

it dont really matter

OpenStudy (anonymous):

Actually, it does.

OpenStudy (anonymous):

Julia can watch 5 hours a week, or less. So which one would that be? \(\geq\), \(\gt\), \(\leq\), or \(\lt\)?

OpenStudy (anonymous):

>

OpenStudy (anonymous):

So Julia can watch more that 5 hours a week? No... Let me describe the symbols...

OpenStudy (anonymous):

\(\geq\) is greater than or equal to. \(\gt\) is greater than

OpenStudy (anonymous):

\(\leq\) is less than or equal to, and \(\lt\) is less than.

OpenStudy (anonymous):

i now what they mean

OpenStudy (anonymous):

So what do you think it is now?

OpenStudy (anonymous):

Hint: Julia can watch 5 hours or less.

OpenStudy (anonymous):

<

OpenStudy (anonymous):

Ugh......................................

OpenStudy (anonymous):

Are you just guessing?

OpenStudy (anonymous):

no <_